OpenStudy (anonymous):

Will Medal And Fan Add. 1 + −2 = Add. −3 + −3 =

OpenStudy (lynfran):

1+(-2) is like 1-2=?? -3+(-3) is like -3-3=??

OpenStudy (anonymous):

yes i think

OpenStudy (lynfran):

ok if you have 1 apple and u take away 2 how much you get left?

OpenStudy (anonymous):

0

OpenStudy (lynfran):

-1

OpenStudy (lynfran):

so, 1-2=-1 ok

OpenStudy (lynfran):

and -3+(-3)=-6 its like the opposite if you had 3+3=6

OpenStudy (lynfran):

u understand?

OpenStudy (anonymous):

sure

OpenStudy (lynfran):

ok